A mole of a newly discovered crystalline substance weighs 14.56 grams and displaces 2.3 ml of water. What is the density of the crystal? Find the atomic radius and density of each atom if it is a simple cubic cell? BCC? FCC? HCP?


1
Expert's answer
2021-11-11T02:34:01-0500

V=2.5mL=2.5cm3 =2.5×10-6 m3

V=L3

2.5×10-6m3 =L3

L=0.09259×10-2m

L=9.259×10-4m


"\therefore" Atomicradius="\frac{1}{2}×length"

"\frac{1}{2}" ×9.259×10-4

=0.00046295m

=462.95μm

Mass=14.56g=0.01456kg

Volume =2.5×10-6 m3

"\therefore" "\rho=\frac{mass}{volume}"


"=0.01456kg\div" 2.5×10-6 m3


=5.824×103kg/m3